Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

a jar has three black marbles and three white marbles.If you draw three marbles replacing each marble before drawing the next one,what is the probability that you will draw at least one black marble?

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

well, we can use complemen probability... 1 - P(3Black) = 1 - (3C3/6C3) = 1 - 1/20 = 19/20

OpenStudy (anonymous):

this is not the answer

OpenStudy (raden):

are there the options

OpenStudy (anonymous):

yup 1/2 3/8 5/8 7/8

OpenStudy (raden):

does even takes marbles one by one or 3 marbles direct ?

OpenStudy (anonymous):

i think one by one

OpenStudy (raden):

oh ya, evens take marble was return... it means P(1B,1W,1W) +P(1B,1B,1W) + P(1B,1B,1B) = 3/6*3/6*3/6 + 3/6*3/6*3/6 + 3/6*3/6*3/6 = 1/8 + 1/8 + 1/8 = 3/8

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!